# Refinement of Monocular Depth Maps via Multi-View Differentiable Rendering

### v1

@@ 3.5 Differentiable Rendering
$\small(\text{I}^{*}_{i},\text{D}^{*}_{i})=\Phi(C_{i},R_{i},t_{i},C_{0},R_{0},t_{0},\mathbf{M}).$ (2)
$\Phi$ projects the current mesh $\mathbf{M}$ with camera parameters $R_{0},t_{0},C_{0}$ to a neighboring view $i$ with $R_{i},t_{i},C_{i}$, creating a pair of synthesized RGB images and depth maps. Specifically, all vertices $p\in\mathbf{M}$ are projected with
$\small p_{i}=C_{i}(R_{i}|t_{i})(R_{0}|t_{0})^{-1}p$ (3)
and then rasterized. $p^{*}_{i}$ is
During optimization, the state of $\mathbf{M}$ continuously changes as described in Sec. 3.6 and 3.7. In the following, we denote the current state of the depth map as $\mathbf{D}_{0}^{*}$. Returning from the meshed representation $\mathbf{M}$ to $\mathbf{D}_{0}^{*}$ is straight forward by evaluating $\Phi$ with $C_{0},R_{0},t_{0}$, which also yields $\mathbf{I_{0}^{*}}$, both of which are rasterized versions of $\mathbf{M}$.
@@ 3.6 Coarse Refinement
In the first optimization stage, the goal is to transfer the initial depth estimation $\mathbf{D}_{0}^{g}$ to more accurate values. Note that in this stage, we do not aim for high precision, but to coarsely align our depth buffer.
To this end, we determine a field that yields an offset $o$ and a scale $s$ for any given image coordinate $(u,v)$, thus mapping depth values $z^{g}\in\mathbf{D}_{0}^{g}$ to refined $z^{c}$:
$\small\begin{array}[]{c}z^{c}=z^{g}(1+s^{c})+o^{c}\quad\quad\forall z^{g}\in\mathbf{D}_{0}^{g},\\ \\ \textrm{with }(o^{c},s^{c})=\Psi(u,v,z^{g}),\end{array}$ (4)
@@
More precisely, the refinement function $\Psi$ is modelled as a shallow multi-layer perceptron (MLP) with positional encoding [48]:
$\small(o^{c},s^{c})=\Psi(\gamma^{m}(u),\gamma^{m}(v),\gamma^{k}(z^{g})),$ (5)
where $\gamma^{m}$ and $\gamma^{k}$ indicate positional encoding functions of degree $m$ and $k$, respectively.
For optimization of the function, we evaluate $\Psi$ directly at available image positions of the sparse point cloud $\mathbf{X}$ and also compare $\mathbf{D}_{0}^{c}$ against $\mathbf{X}$ and use the regularizer as described in Sec. 3.8 using gradient descent. Evaluating the losses of $\mathbf{D}_{0}^{c}$ will prevent overfitting of $\Psi$ solely to the available points. The weights of $\Psi$ are initialized very small, thus $o^{c}$ and $s^{c}$ will be small and thus $z^{c}$ close to $z^{g}$.
@@ 3.7 Local Refinement
The second stage of our refinement aims to finely tune $\mathbf{D}_{0}^{c}$, arriving at $\mathbf{D}_{0}$. Therefore, we jointly optimize depth values $z$ with vertex image coordinates $u,v$ and interject mesh decimation steps in between. The local refinement step aligns the surface mesh with edges in the image and adjusts depth values to finely describe the surface.
Depth Refinement.
To efficiently optimize depths, we introduce an inverse scaling factor $f_{z}$ per vertex which serves as the optimizable parameter.
$\small z=\frac{z^{c}}{f_{z}}$ (6)
Note that local depth refinement is not optimized during the coarse refinement, but rather started afterwards.
Coordinate Refinement.
We further introduce the optimizable image space offsets $\Delta u,\Delta v$ per vertex, thus allowing for a piece-wise linear approximation of silhouettes. $\Delta u,\Delta v$ are applied to the vertices with
$\small(u,v)=(u^{g}+\Delta u,v^{g}+\Delta v).$ (7)
$\Delta u,\Delta v$ are limited by $\frac{d}{2}$ to ensure that the mesh vertices won’t intersect. Note that in contrast to $f_{z}$, $\Delta u,\Delta v$ are optimized already during the coarse refinement.
As in the previous refinement, $\Delta u,\Delta v$ and $f_{z}$ are optimized with gradient descent as $\Phi$ is evaluated repeatedly with $(C_{i},R_{i},t_{i})$ and optimized according to the losses described in the following section.
@@ 3.8 Losses & Regularizers
In this section, we describe the two losses and three regularizers that are used to control and stabilize our optimization.
@@
${\mathcal{R}_{\textit{s}}}$ is a simple regularizer that suppresses noisy results:
$\small{\mathcal{R}_{\textit{s}}}(\mathbf{D}_{0}^{*})=||\mathbf{D}_{0}^{*}-g(\mathbf{D}_{0}^{*})||_{2},$ (8)
where $g$ is a Gaussian blur with window size 5.
The regularizers $\mathcal{R}_{\textit{p}}$ and $\mathcal{R}_{\textit{e}}$ have the goal to maintain structural information from the mono depth image $\mathbf{D}_{0}^{r}$. To this end, we calculate edge images $\mathbf{E}_{0}^{r}$ and $\mathbf{E}_{0}^{*}$ from $\mathbf{D}_{0}^{r}$ and $\mathbf{D}_{0}^{*}$ by utilizing a Sobel filter-based edge detection. The edge images $\mathbf{E}$ have values $\in[-1,1]$, with values below 0 indicating edges and discontinuities and values above 0 indicating surfaces. Following, we define the edge regularizer $\mathcal{R}_{\textit{e}}$ as
$\small\begin{array}[]{c}\mathcal{R}_{\textit{e}}(\mathbf{E}_{0}^{*},\mathbf{E}_{0}^{r})=||\max(0,\mathbf{E}_{0}^{*})-\max(0,\mathbf{E}_{0}^{r})||_{2}+\frac{1}{|\mathbf{M}|}\sum\limits_{j=1}^{|\mathbf{M}|}\mathbf{e}_{u^{\prime},v^{\prime}}\end{array}$ (9)
Inspired by Poisson blending [37], we define $\mathcal{R}_{\textit{p}}$ as:
$\small\begin{array}[]{c}\mathcal{R}_{\textit{p}}(\mathbf{D}_{0}^{*},\mathbf{D}_{0}^{r})=||\max(0,\mathbf{E}_{0}^{r})(\nabla_{u}\tilde{\mathbf{D}}_{0}^{*}-\nabla_{u}\mathbf{D}_{0}^{r})||_{2}+\\ \\ ||\max(0,\mathbf{E}_{0}^{r})(\nabla_{v}\tilde{\mathbf{D}}_{0}^{*}-\nabla_{v}\mathbf{D}_{0}^{r})||_{2}.\end{array}$ (10)
@@
$\small\mathcal{L}_{\textit{photo}}(\mathbf{I}^{*}_{i},\mathbf{I}_{i})=||\max(0,\mathbf{E}_{i}^{*})(\mathbf{I}^{*}_{i}-\mathbf{I}_{i})||_{2}.$ (13)
@@
$\max(0,\mathbf{E}_{i}^{*})$ indicates the projected depth map $\mathbf{E}_{r}^{*})$ will mask out regions pf $\mathbf{I}_{i}$ which were occluded in $\mathbf{I}_{0}$.
@@
Our total evaluation loss combines the introduced objective functions using a weighted sum of each of the elements, see supplemental material.
4 Depth Refinement with GPU Acceleration
Using a GPU hardware accelerated framework for our depth refinement is preferable due to implementation simplicity and optimization speed. This however is challenging as common hardware implementations (e.g. NVDiffRast [26]) follow hardware specifications and use a non-linear, reciprocal depth formulation, normalized between $[-1,1]$ after screen projection. Additionally, this formulation uses a near plane $d_{\textit{near}}$ and far plane $d_{\textit{far}}$ which restricts the minimal and maximal length of the camera ray of the principal point resulting in a $4\times 4$ matrix $P$.

### v2

@@ 3.4 Differentiable Rendering
$(\mathbf{I}^{*}_{i},\mathbf{D}^{*}_{i})=\Phi(\textbf{C}_{i},\textbf{R}_{i},\textbf{t}_{i},\textbf{C}_{0},\textbf{R}_{0},\textbf{t}_{0},\mathbf{M}).$ (2)
$\Phi$ projects the current mesh $\mathbf{M}$ with camera parameters $\textbf{R}_{0},\textbf{t}_{0},\textbf{C}_{0}$ to a neighboring view $i$ with $\textbf{R}_{i},\textbf{t}_{i},\textbf{C}_{i}$, creating a pair of synthesized RGB images and depth maps. Specifically, all vertices $\textbf{p}\in\mathbf{M}$ (see Eq. 1) are projected with
$\textbf{p}^{*}_{i}=\textbf{C}_{i}[\textbf{R}_{i}|\textbf{t}_{i}][\textbf{R}_{0}|\textbf{t}_{0}]^{-1}\textbf{p}$ (3)
and then rasterized. During optimization, the state of $\mathbf{M}$ continuously changes as described in Sec. 3.5 and 3.6. In the following, we denote the current state of the depth map as $\mathbf{D}_{0}^{*}$. Returning from the meshed representation $\mathbf{M}$ to $\mathbf{D}_{0}^{*}$ is straight forward by evaluating $\Phi$ with $i=0$, thus $\textbf{C}_{0},\textbf{R}_{0},\textbf{t}_{0}$, which also yields $\mathbf{I}_{0}^{*}$, both of which are rasterized versions of $\mathbf{M}$.
@@ 3.5 Coarse Refinement
In the first optimization stage, the goal is to transfer the initial depth estimation $\mathbf{D}_{0}^{g}$ to more accurate values. Note that in this stage, we do not aim for highest precision, but mainly to coarsely align our depth buffer to the sparse SfM point cloud having only a few degrees of freedom.
To this end, we determine a neural field that yields an offset $o$ and a scale $s$ for any given image coordinate $(u,v)$, thus mapping depth values $z^{g}\in\mathbf{D}_{0}^{g}$ to refined $z^{c}$:
$\small\begin{array}[]{c}z^{c}=z^{g}(1+s)+o\quad\quad\forall z^{g}\in\mathbf{D}_{0}^{g},\\ \\ \textrm{with }(o,s)=\Psi(u,v,z^{g}),\end{array}$ (4)
@@
Specifically, $\Psi$ is modeled as a shallow multi-layer perceptron (MLP):
$\small(o,s)=\Psi(\gamma^{m}(u),\gamma^{m}(v),\gamma^{k}(z^{g})),$ (5)
where $\gamma^{m}$ and $\gamma^{k}$ indicate positional encoding [55] functions of degree $m$ and $k$, respectively.
For optimization of the function, we evaluate $\Psi$ directly at available image positions of the sparse point cloud $\mathbf{X}$ and also compare $\mathbf{D}_{0}^{c}$ against $\mathbf{X}$ and use the regularizer as described in Sec. 3.7 using gradient descent. Evaluating the losses of $\mathbf{D}_{0}^{c}$ will prevent overfitting of $\Psi$ solely to the available points. The weights of $\Psi$ are initialized very small (see the Appendix B for details), thus $o^{c}$ and $s^{c}$ will be small, and thus $z^{c}$ close to $z^{g}$.
@@ 3.6 Local Refinement
The last refinement stage aims to finetune $\mathbf{D}_{0}^{c}$, arriving at $\mathbf{D}_{0}$. Therefore, we jointly optimize depth values $z$ with vertex image coordinates $u,v$ and interject mesh decimation steps in between. The local refinement step aligns the triangle edges with silhouettes in the image and adjusts depth values to accurately trace the surface.
We introduce the per-vertex inverse scaling factor $f_{z}$ and image space offsets $\Delta u,\Delta v$, thus allowing for a piece-wise linear approximation of silhouettes, as optimizable parameters:
[…]
$\Delta u,\Delta v$ are limited by $\frac{d}{2}$ to ensure that the mesh vertices won’t intersect. (Note that in practice, we enable optimization of $\Delta u,\Delta v$ already during the coarse refinement.)
As in the previous refinement, $\Delta u,\Delta v$ and $f_{z}$ are optimized with gradient descent as $\Phi$ is evaluated repeatedly with $(\mathbf{C}_{i},\mathbf{R}_{i},\mathbf{t}_{i})$ and optimized according to the losses described in Sec. 3.7.
For challenging datasets with varying lighting conditions, we also introduce a tonemapping module that optimizes per image exposure and response parameters. For details, we refer the reader to Appendix F.
@@ 3.7 Losses & Regularizers
We outline losses and regularizers used for the optimization. Exact formulas and setups are in Appendix A.
Geometric Loss.
We define $\mathcal{L}_{\textit{geo}}$ as the relative Huber loss [21] between the current depth buffer and the depth values of the projected sparse point cloud. To get the correspondence between the $\mathbf{D}^{*}_{0}$ and $\mathbf{X}$, we calculate $\mathbf{R}_{0}|\mathbf{t}_{0}]X$ to get the linear $z^{x}$, and $\mathbf{C}_{0}[\mathbf{R}_{0}|\mathbf{t}_{0}]X$ to get the pixel coordinates per point. Note that we use the point visibility information that COLMAP provides, so we only consider points that are actually visible in $I_{0}$. We use $\mathcal{L}_{\textit{geo}}$ during coarse and local refinement directly on $\mathbf{D}^{*}_{0}$. During coarse refinement, we employ $\mathcal{L}_{\textit{geo}}$ on $z^{c}$ (see Eq. 4) with $\mathbf{X}$.
Monocular Sobel Filtered Projected
Depth Map $\mathbf{D}_{0}^{r}$ Edge Map $\mathbf{E}_{0}$ Edge Map $\mathbf{E}_{i}^{*}$
Figure 6: To ensure proper visibility masking in nearby-view rendering, we project the edge map $\mathbf{E}_{0}$ onto the rendered view to weight color loss (white fully weighted, black not weighted).
Photometric Loss.
We use a standard mean squared error formulation for our photometric consistency loss $\mathcal{L}_{\textit{photo}}$ between rendered $\mathbf{I}^{*}_{i}$ and captured images $\mathbf{I}_{i}$. However, $\mathbf{D}_{0}$ cannot represent what is behind visible objects in the scene, so as we render $\mathbf{I}^{*}_{i}$, regions that were occluded in $\mathbf{I}_{0}$ are revealed. To prevent false comparisons, such regions are masked in the loss. To this end, we rely on $\mathbf{D}_{0}^{r}$ and calculate its edge image $\mathbf{E}_{0}$ by utilizing a Sobel filter-based edge detection. We use the the projected edge map $\mathbf{E}_{i}^{*}$ to mask out the loss in occluded regions, which can be seen in Fig. 6.
Optimizing for photometric consistency presents a non-convex challenge with a chance of becoming trapped in local minima if the optimum is too distant. Thus, we opted to only apply this loss during local refinement when our current depth estimate $\mathbf{D}^{*}_{0}$ is closer to the optimum.
@@
To stabilize the optimization, we apply different regularizers, which carefully balance the contributions of dense photometric consistency, sparse geometric cues and the initial estimation. Their general idea is to incorporate and reuse the monocular depth image $\mathbf{D}_{0}^{r}$ as prior knowledge about structural information and scene layout and apply that to the meshed and rendered version $\mathbf{D}_{0}^{*}$.
The edge targeting regularizer $\mathcal{R}_{\textit{e}}(\mathbf{E}_{0}^{*},\mathbf{E}_{0})$, with $\mathbf{E}_{0}^{*}$ being the edge image of $\mathbf{D}_{0}^{*}$, encourages that silhouettes of the rendered depth map $\mathbf{D}_{0}^{*}$ and estimated depth map $\mathbf{D}_{0}^{r}$ should align by calculating their mean squared error and an additional term that moves $(u,v)$ on to the edges by minimizing the values sampled from $\mathbf{E}_{0}$.
A Poisson blending-inspired regularizer $\mathcal{R}_{\textit{p}}$ is proposed for preserving the local structure from $\mathbf{D}_{0}^{r}$ to $\mathbf{D}_{0}^{*}$ by assessing their image space gradients. Again we mask out silhouettes using $\mathbf{E}_{0}^{r}$ to neglect discontinuities, which are prone to be inaccurate. An alternative to reusing the depth map gradients is to regularize on available normals. Notably, $\mathcal{R}_{\textit{p}}$ performed on par with using a SotA normal estimator [1] (see App. F) and demonstrating that our regularizer should be well applicable also outside of the context of this work. Our total evaluation loss combines the introduced objective functions using a weighted sum of each of the elements, see Appendix B.
4 Results
4.1 Setup & Datasets
We evaluate our method on the synthetic Replica dataset [53, 45] and the real-world dataset ScanNet++ [73]. Both offer dense metric scale ground truth, thus allowing for an isolated evaluation of individual depth maps.
For Replica, we use color and ground truth depth maps from rendered meshes, from 2 room and 2 office scenes randomly selected from the subset prepared by Rosinol et al. [45] and optimizing over 64 views per scene. In ScanNet++, ground truth depth maps come from meshes, and the sparse point clouds and poses are computed by COLMAP [47]. For results covered in Sec. 4.2, we used 10 ScanNet++ validation scenes, each with 10 depth maps, totaling 100. Scenes where chosen to have little blurry images, and only consecutive frames. More details regarding the dataset preparation are available in Appendix D.
We evaluate standard error metrics: root mean squared error (RMSE), mean absolute error (MAE), L1-rel [57], L1-inv [57], and $\delta_{1}$ [69]. We also assess depth sample accuracy ($|z-z_{GT}|<\tau$) at various thresholds $\tau$. Additionally, we measure the density of depth maps by counting the number of valid pixels. Values in the top 20 percent are indicated in green, the greener the better.
Notably, we used the exact same hyperparameters for all scenes and datasets, highlighting the stability of our pipeline. This set of parameters was determined by an extensive set of experiments on withheld ScanNet++ scenes. Ablation studies regarding the size of the involved mini MLP during coarse refinement, the mesh decimation ratios $d$ and $r$, the tonemapper, different monocular depth estimators (generally, we used Marigold [25]), and multi depthmap refinement can be found in Appendix B and C. Additionally, we evaluate various loss configurations, even using additional normal estimators, which underlines the flexibility of our method as information from different sources can be trivially fused by adding a loss or adapting the initialization of $\mathbf{D}_{0}^{*}$.
$\downarrow$ RMSE MAE L1-inv $\uparrow$ Acc. Acc. Acc. Cumul.
$\tau$=0.01 m $\tau$=0.05 m $\tau$=0.10 m time [s]
Global estimation 0.29 0.21 0.17 0.06 0.27 0.47
(least squares)
After step:
Init. & our global est. ($\mathbf{D}^{g}_{0}$) 0.20 0.14 0.11 0.08 0.32 0.55 11.2
Meshing ($\mathbf{M}$) 0.20 0.14 0.11 0.08 0.32 0.55 13.3
Coarse alignment ($\mathbf{D}^{c}_{0}$) 0.22 0.12 0.12 0.12 0.44 0.66 128
Local refinement ($\mathbf{D}_{0}$) 0.17 0.09 0.07 0.22 0.58 0.76 437
Table 1: Intermediate results. On the ScanNet++ test data, each step of our method significantly improves intermediate outcomes, outperforming the least squares baseline [25]. Notably, meshing the depth map preserves quality from $\mathbf{D}^{g}_{0}$. Time measured on a Nvidia RTX3090 with OpenGL backend.
See Tab. 1 for a detailed summary of the efficiency each involved step. Each step of our method notably enhances results, surpassing the least squares baseline [25].
While this section mainly focuses on per depth map quality, we further evaluated scene global reconstruction quality on a subset of the Tanks and Temples [28] dataset, showing similar results compared to COLMAP.
4.2 Comparison against Prior Arts
[…]
Method $\uparrow$ Valid $\downarrow$ RMSE MAE L1-rel $\uparrow$ $\delta_{1}$ Acc. Acc. Acc.
Samples 0.01 m 0.05 m 0.10 m
MVSFormer (0) 1.00 0.80 0.31 0.13 0.88 0.21 0.58 0.73
MVSFormer++ (0) 1.00 0.56 0.27 0.10 0.89 0.23 0.61 0.76
MVSFormer (0.25) 0.97 0.75 0.27 0.11 0.90 0.21 0.58 0.73
MVSFormer++ (0.25) 0.95 0.42 0.17 0.06 0.93 0.23 0.61 0.76
MVSFormer (0.5) 0.87 0.46 0.14 0.06 0.96 0.21 0.57 0.71
MVSFormer++ (0.5) 0.85 0.19 0.07 0.01 0.98 0.22 0.59 0.73
Ours (+) 0.99 0.08 0.04 0.02 1.00 0.37 0.81 0.92
Ours (-) 0.99 0.09 0.05 0.02 0.99 0.26 0.69 0.87
Table 2: Results on the Replica dataset. In brackets are confidence thresholds for MVSNet-based methods and sparse point cloud accuracy with good (+) and poor (-) quality for our method.
For the synthetic Replica dataset, sparse point clouds are generated from ground truth data with added realistic levels of noise and outliers. Our method excelled the dataset, as noted in Tab. 2, regardless of point cloud quality. Here, ’Ours (+)’ denotes high-quality point clouds (2% noise + 2% outliers). In contrast, ’Ours (-)’indicates use of very low-quality point clouds (5-10% noise + 5-10% outliers) leading to artifacts in extreme cases (see Appendix D).
Method $\uparrow$ Valid $\downarrow$ RMSE MAE L1-rel $\uparrow$ $\delta_{1}$ Acc. Acc. Acc.
Samples 0.01 m 0.05 m 0.10 m
DepthAnything2
(metric) 100% 1.91 1.79 1.51 0.00 0.00 0.00 0.00
COLMAP 69% 0.60 0.30 0.24 0.73 0.17 0.35 0.41
MASt3R 98% 0.19 0.16 0.16 0.89 0.07 0.30 0.51
MVSFormer0 100% 1.08 0.47 0.41 0.68 0.16 0.40 0.52
MVSFormer++0 100% 0.67 0.28 0.24 0.79 0.22 0.53 0.65
GeoMVSNet0 100% 1.19 0.61 0.77 0.53 0.05 0.18 0.28
MVSFormer0.25 90% 1.09 0.46 0.38 0.72 0.16 0.40 0.51
MVSFormer++0.25 88% 0.61 0.22 0.19 0.86 0.22 0.52 0.63
GeoMVSNet0.25 100% 1.19 0.61 0.77 0.53 0.05 0.18 0.28
MVSFormer0.5 47% 0.52 0.16 0.10 0.93 0.14 0.30 0.36
MVSFormer++0.5 62% 0.26 0.10 0.07 0.96 0.19 0.43 0.50
GeoMVSNet0.5 23% 1.45 0.90 1.11 0.38 0.00 0.02 0.04
Ours 99% 0.17 0.090 0.07 0.93 0.22 0.58 0.76
Table 3: Results on the ScanNet++ dataset [73]. Numbers in subscript indicate the confidence threshold for MVSNet methods.
Quantitative results for ScanNet++ are presented in Tab. 3. We achieved top or on-par results in RMSE, completeness (valid samples), MAE, L1-rel, and accuracy at $\tau=0.05$ and $\tau=0.1$. MVSFormer++ with a certainty threshold of 0.5 slightly outperformed us in L1-inv, $\delta_{1}$, and accuracy at $\tau=0.01$. However, their results only included 62% of values, compared to our over 99.5% of samples.
The qualitative results and ScanNet++ are shown in Fig. 7. Our outputs are clean, dense, and plausible. However, some artifacts appear on the glossy table in Scene 3. Sometimes, our results surpass the ground truth, as in the crop of Scene 4 or the table leg in Scene 5, right crop. MVSNet-like results are sparse with noisy walls. DepthAnything2’s metric version offers promising details, though its global scale is inaccurate.
Figure 7: Qualitative results on the ScanNet++ dataset, with MVSFormers’ certainty threshold at their default of 0.5. Green borders indicate image crops, while red highlights error maps; brighter areas imply greater errors, and grey indicates missing samples.
5 Limitations
While our method shows strong performance, it still has limitations. Photometric consistency proves useful, but fails in cases with specular materials (see e.g. Fig. 7) which our method does not account for. This could be addressed with uncertainty estimation, anchoring optimization with high-certainty depth estimates.
Ultimately, our approach relies on the initial input. Outliers in the sparse point cloud can result in spurious artifacts and the scene topology, in form of an edge image, has to be derived from the monocular depth map. The edge image is crucial for accurately masking occluded areas, influencing effectiveness of the losses.
6 Conclusion
In conclusion, our approach demonstrates that monocular depth maps can be significantly improved by incorporating photometric consistency across multiple views, leading to enhanced depth estimation. By optimizing a meshed version of the monocular depth map with a differentiable renderer, we directly leverage the strong initial estimates and achieve stable optimization, which refines scaling and corrects errors. In this context, our Poisson and edge regularizers have proven to be particularly effective, ensuring robust and reliable depth map refinement. Our simple yet effective pipeline consistently produces accurate and dense depth maps, which is highly desirable for downstream tasks.
7 Acknowledgments
Appendix A Details on Losses & Regularizers
In this section, we describe the geometric, photometric and dense geometric consistency losses and regularizers that are used to control and stabilize our optimization in more detail.
@@ Appendix A Details on Losses & Regularizers
[…]
@@
$\max(0,\mathbf{E}_{i}^{*})$ indicates the projected edge map $\mathbf{E}_{r}^{*}$ will mask out regions of $\mathbf{I}_{i}$ which were occluded in $\mathbf{I}_{0}$.
Regularizers.
The regularizers $\mathcal{R}_{\textit{p}}$ and $\mathcal{R}_{\textit{e}}$ have the goal to maintain structural information from the mono depth image $\mathbf{D}_{0}^{r}$. To this end, we calculate edge images $\mathbf{E}_{0}^{r}$ and $\mathbf{E}_{0}^{*}$ from $\mathbf{D}_{0}^{r}$ and $\mathbf{D}_{0}^{*}$ by utilizing a Sobel filter-based edge detection. The edge images $\mathbf{E}$ have values $\in[-1,1]$, with values below 0 indicating edges and discontinuities and values above 0 indicating surfaces.
Following, we define the edge regularizer $\mathcal{R}_{\textit{e}}$ as
[…]
Inspired by Poisson blending [43], we define $\mathcal{R}_{\textit{p}}$ as:
[…]
@@
Our total evaluation loss combines the introduced objective functions using a weighted sum of each of the elements, see Tab. 4.
Loss Regularizer
$\mathcal{L}_{\textit{photo}}$ 1.0 $\mathcal{R}_{\textit{p}}$ 400.0
$\mathcal{L}_{\textit{geo}}$ 0.1 $\mathcal{R}_{\textit{e}}$ 0.1
($\mathcal{L}_{\textit{photo}}^{\text{MV}}$ 1.0) ($\mathcal{R}_{\textit{s}}$} 0.01)
($\mathcal{L}_{\textit{gc}}$} 0.1) ($\mathcal{R}_{n}$ 0.1)
Table 4: Used weights for computing the total optimization loss. Brackets indicate usage only in ablation studies.
Appendix B Training Details
For our implementation, we use the hardware accelerated framework NVDiffRast [30] with the adjusted formulation outlined in Sec. C. For initial depth estimations, we use Marigold [25], if not stated otherwise, and employ COLMAP [47, 48] for Structure-from-Motion and image undistortion with configurations as recommended by the used dataset [73]. In Fig. 8, we depict sparse pointclouds of representative views, showcasing sparseness especially on white walls. For coarse refinement, $\Psi$ is a shallow MLP with degrees of positional encoding being $m=3$ and $k=5$. Two hidden layers of width 16, initialized with a normal distribution with $\sigma=0.1$. This comparatively small initialization ensures $o_{g}$ and $s_{g}$ being very small initially, thus, starting close to the initial depth estimation. During the coarse refinement phase, we used a learning rate of $0.001$ and stop after 400 iterations with batch sizes of 16. The current states of $o_{g}$ and $s_{g}$ are baked into $\mathbf{M}$’s $z$. We stop optimization after another 1100 iterations of local refinement with learning rates of $0.0005$. The number of iterations were found empirically based on our ScanNet++ validation set. In total, the process takes around 8 min per depth map (including output writing and logging) on a NVIDIA RTX 3090 using the CUDA backend. We see potential for performance improvements by e.g. replacing the MLP of the coarse refinement by an explicit data structure that can be optimized faster. We used the Adam [27] optimizer with default parameters as proposed by the PyTorch framework [52].
Appendix C Implementation with GPU Acceleration
Using a GPU hardware accelerated framework (we use NVDiffRast [30]) is preferable due to implementation simplicity and optimization speed. To leverage hardware acceleration, our implementation must follow specifications as defined by the OpenGL convention [3]. As a consequence, we must fulfill the following conditions: 1) The intrinsics matrix $C$ must be reformulated as a $4\times 4$ projection matrix [3], which limits the visible depth range by to the range $z_{\textit{near}},z_{\textit{far}}$] and normalizes image space positions to $1,1]^{3}$, referred to as clip space [3]. 2) Consequently, we also need to take care of edge case handling when depth values approach these near and far values. 3) As the coordinate refinement of $(u,v)$ also has to be performed in clip space, we additionally need to store the depth of $\mathbf{M}$ in non-linear reciprocal space. Optimizing in reciprocal space would be very instable, so we perform depth refinement distinctly using a linear scaling, as described in Sec. 3.6.
In the following, we describe the adaptions made to comply with the mentioned requirement in more detail.
Projection Matrices.